the diameter on a wheel on a car is 1.5 feet. how many revolutions must the wheel make so that it travels 37.68 feet. use pi=3.14

Answer by bucky(2189)

The circumference of the wheel is found using the formula:
.
C = pi*D
.
where C is the circumference and D is the diameter. So for this problem:
.
C = pi*1.5 = 3.14*1.5 = 4.71 feet
.
Each turn of the wheel moves the car forward a distance equal to the circumference of the wheel
.
To move the car forward 37.68 feet the number of rotations (revolutions) the wheel must make
is found by dividing 4.71 into 37.68 ....
.
Rotations = 37.68/4.71 = 8
.
So, if the wheel rotates 8 times, the car will move forward 37.68 feet.
.
Hope this helps you to understand the problem.
.